Axolotl:

The diagonals of parallelogram LMNO intersect at point P. If MP = 2x − 1 and OP = 3x − 4, what is MP? −1 1 3 5

Axolotl:

Oh hi Astrid! We are both axolotls lol

Astrid1:

We're given: MP=2x-1 OP=3x-1 Equate The Diagonals: MP=OP 2x-1=3x-4 Collect Like Terms (x) 2x-3x=-4+1 -x=-3 Make (x) the subject x=-3/1 x=3 MP= 2x-1=2(3)-1=6-1=5 So your answer should be option d, 5. xd hello

Axolotl:

@astrid1 wrote:
We're given: MP=2x-1 OP=3x-1 Equate The Diagonals: MP=OP 2x-1=3x-4 Collect Like Terms (x) 2x-3x=-4+1 -x=-3 Make (x) the subject x=-3/1 x=3 MP= 2x-1=2(3)-1=6-1=5 So your answer should be option d, 5. xd hello
THANK YOU SO MUCH! I didn't really know how to do this ;-;

Astrid1:

Ofc!! I hope I helped ;c

Astrid1:

Haven't done problems like this in a little bit

Axolotl:

@astrid1 wrote:
Haven't done problems like this in a little bit
Yes it helped me thank you <3 (you are now one of my favorites! It is not just because you are also a Axolotl now lol)

Axolotl:

@astrid1 wrote:
Ofc!! I hope I helped ;c
I really appreciate that you took time to answer my question and that you helped me by doing it step by step

jhonyy9:

@axolotl wrote:
@astrid1 wrote:
We're given: MP=2x-1 OP=3x-1 Equate The Diagonals: MP=OP 2x-1=3x-4 Collect Like Terms (x) 2x-3x=-4+1 -x=-3 Make (x) the subject x=-3/1 x=3 Attention !!! x = -3/(-1) not +1 for x= -3/1 you get x= -3 not 3 MP= 2x-1=2(3)-1=6-1=5 So your answer should be option d, 5. xd hello
THANK YOU SO MUCH! I didn't really know how to do this ;-;

Axolotl:

@jhonyy9 wrote:
@axolotl wrote:
@astrid1 wrote:
We're given: MP=2x-1 OP=3x-1 Equate The Diagonals: MP=OP 2x-1=3x-4 Collect Like Terms (x) 2x-3x=-4+1 -x=-3 Make (x) the subject x=-3/1 x=3 Attention !!! x = -3/(-1) not +1 for x= -3/1 you get x= -3 not 3 MP= 2x-1=2(3)-1=6-1=5 So your answer should be option d, 5. xd hello
THANK YOU SO MUCH! I didn't really know how to do this ;-;
Why did you repost the messages? You didn't say anything.

jhonyy9:

@axolotl wrote:
@jhonyy9 wrote:
@axolotl wrote:
@astrid1 wrote:
We're given: MP=2x-1 OP=3x-1 Equate The Diagonals: MP=OP 2x-1=3x-4 Collect Like Terms (x) 2x-3x=-4+1 -x=-3 Make (x) the subject x=-3/1 x=3 Attention !!! x = -3/(-1) not +1 for x= -3/1 you get x= -3 not 3 MP= 2x-1=2(3)-1=6-1=5 So your answer should be option d, 5. xd hello
THANK YOU SO MUCH! I didn't really know how to do this ;-;
Why did you repost the messages? You didn't say anything.
bc. there is wrong what sou have wrote please correct it x = -3/1 = -3 not +3

Axolotl:

@jhonyy9 wrote:
@axolotl wrote:
@jhonyy9 wrote:
@axolotl wrote:
@astrid1 wrote:
We're given: MP=2x-1 OP=3x-1 Equate The Diagonals: MP=OP 2x-1=3x-4 Collect Like Terms (x) 2x-3x=-4+1 -x=-3 Make (x) the subject x=-3/1 x=3 Attention !!! x = -3/(-1) not +1 for x= -3/1 you get x= -3 not 3 MP= 2x-1=2(3)-1=6-1=5 So your answer should be option d, 5. xd hello
THANK YOU SO MUCH! I didn't really know how to do this ;-;
Why did you repost the messages? You didn't say anything.
bc. there is wrong what sou have wrote please correct it x = -3/1 = -3 not +3
What would the answer be then? I'm trying to figure it out.

jhonyy9:



Axolotl:

@jhonyy9 wrote:
Created with RaphaëlLMNOPMP = 2x -1OP = 3x -4MP = OPReply Using Drawing
Ok. So I'm trying to do the math. 2x - 1 = 3x - 4

jhonyy9:

from this equality you need calcule the x value x = ?

Axolotl:

@jhonyy9 wrote:
from this equality you need calcule the x value x = ?
I know! That's what I'm struggling to figure out :(

Axolotl:

@axolotl wrote:
@jhonyy9 wrote:
from this equality you need calcule the x value x = ?
I know! That's what I'm struggling to figure out :(
ACKKKKKKKKK this is why I hate Geometry

jhonyy9:

2x -1 = 3x -4 subtract from both sides 2x so what you get ?

Axolotl:

@jhonyy9 wrote:
2x -1 = 3x -4 subtract from both sides 2x so what you get ?
2x -1 = 3x - 4 -2x -2x -1 = x - 4

jhonyy9:

yes and now x = ?

Axolotl:

WAIT I THINK I MIGHT HAVE IT

Axolotl:

YASSSS

jhonyy9:

add to both sides 4

jhonyy9:

@axolotl wrote:
@jhonyy9 wrote:
2x -1 = 3x -4 subtract from both sides 2x so what you get ?
2x -1 = 3x - 4 -2x -2x -1 = x - 4
and now add to both sides 4

Axolotl:

@jhonyy9 wrote:
add to both sides 4
2x -1 = 3x - 4 -2x -2x -1 = x - 4 +4 + 4 3 = x Am I correct?

jhonyy9:

yes of cours x = 3 but you need get the MP = ? but there is MP = 2x-1 so now just substitute the 3 in please of x and will get the MP

Axolotl:

oh ok

jhonyy9:

MP = ?

Axolotl:

2(3) − 1 6 - 1 = 5

jhonyy9:

yes perfect MP = 5 but now than you wan check your answer is correct we know that MP = OP OP = 3x -4 for x = 3 you need get OP = MP = 5 do you think this is true ?

jhonyy9:

OP = 3x -4 substitute the 3 in place of x and you need get 5

Axolotl:

3(3) - 4 9 - 4 5

jhonyy9:

perfect so do you understand it all ?

Axolotl:

@jhonyy9 wrote:
perfect so do you understand it all ?
YES! Thank you so much! I really really appreciate this!

jhonyy9:

np anytime just tag me when i m here online hnd cya

Axolotl:

@jhonyy9 wrote:
np anytime just tag me when i m here online hnd cya
Bye thank you so so much! Both you and @astrid1 were a big help! Thank you both! No one usually helps me with my work so I'm so happy!